OpenStudy (anonymous):

Write the equation of the quadratic function with roots -1 and 1 and a vertex at (0, -3).

OpenStudy (loser66):

where are you stuck?

OpenStudy (anonymous):

I am just not sure of the format in which to write this type of equation. Again I was not thoroughly introduced to the unit so I am having trouble with the basics

OpenStudy (loser66):

quadratic means the exponent of x is 2, right?

OpenStudy (loser66):

The vertex form of a quadratic is y = a(x-h)^2 +k, right?

OpenStudy (loser66):

vertex (h,k) = (0,-3) , hence y = ax^2 -3 now, it has roots, that is y =0 at those roots. Just pick x =-1, put it into the function to get a 0=a(-1)^2 -3--> a =3 now, you have all information, just plug in y = 3x^2 -3

OpenStudy (anonymous):

awesome, thanks!